# Prime and Reach: Synthesising Body Motion for Gaze-Primed Object Reach

### v1

@@ 5.3 Baselines
To assess the ability of generated motion to replicate humans’ P&R behaviour, we benchmark five methods on our curated datasets: one naive baseline and three previous works (one from the text-to-motion generation and two from location-conditioned human motion generation):
  • Static is a naive baseline that uses the average full-body pose of training data and keeps it static for the entire motion sequence. It showcases the difficulty of the dataset.
  • MDM [tevet2023mdm]. We evaluate the checkpoint trained on HumanML3D [Guo_2022_CVPR] vs our pre-training on Nymeria [ma2024nymeria]. We also fine-tune this model, pre-trained on Nymeria, on each dataset using only text conditioning.
  • GMD [karunratanakul2023gmd], a guided motion diffusion trained on HumanML3D based on UNet architecture, is a two-stage process designed for controllable human motion synthesis with text. The first stage generates a root trajectory that guides full-body motion generation in the second stage.
  • WANDR [diomataris2024wandr] is a data-driven, autoregressive method that uses a conditional Variational Auto-Encoder (c-VAE) to generate motion frame-by-frame, trained on the combined dataset of AMASS [mahmood2019amass] and CIRCLE [araujo2023circle]. The model is guided by intention features that encode the goal’s position and the time remaining to reach it.
@@ 5.4 Results
In Tab. 2, we compare our proposed P&R diffusion model against baselines on five different datasets. The naive static baseline performs poorly on all the metrics for all datasets. Its poor prime success ($20\%$ on average) and reach success ($12\%$ on average) highlight the difficulty of the task on our curated datasets.
@@
Text-conditioned baselines rely solely on the knowledge of the pick/put action to generate motion and have no information about the target location to prime or reach. We report the performance of three variants of MDM [tevet2023mdm]. These baselines perform poorly on all the metrics as they lacks sufficient guidance. Finetuning MDM on the target datasets helps generate better prime and reach motions with a maximum gain of $24.7\%$ in prime success and $17.1\%$ in reach success on HD-EPIC. The fine-tuned MDM only marginally improves over the static baseline (average of $20.0\%$ prime success and $13.1\%$ reach success), showing that text is not a sufficient condition for prime and reach.
[…]
@@
Figure 5: Qualitative results on 3 datasets: Ground truth sequence in light green, goal-pose conditioned prediction in translucent yellow, and target location conditioned generation in brown. We show the pose at the initial, prime, and reach timesteps. Prime direction for both ground truth and predictions are shown using arrows, and target object location is shown in sphere.
Conditioning with Object Location. WANDR [diomataris2024wandr] achieves strong success scores over all datasets, as that the model’s reward focuses on the reaching of goal. However, it underperforms on all other metrics, including the metrics that evaluate the motion between the initial state and target, and it is very poor at priming, with an average of only $16\%$ over all datasets. P&R trained using initial state and object location as condition achieves significant improvements in prime success with gains ranging from $11.9\%$ on ADT to $48.8\%$ on HD-EPIC. It reaches or outperforms WANDR in reach success on three out of the five datasets. It also outperforms WANDR by a big margin on all other metrics.
Importantly, when comparing the object location condition with the goal pose, we note that metrics that evaluate the goal location (Goal MPJPE and Loc Error) surely benefit from the guidance on where the person is exactly standing at the end. However, when focusing on the prime and reach success, object location can outperform the stronger goal pose on MoGaze and GIMO. We find that due to limited training data, performance on small-scaled ADT and GIMO are evidently worse.
@@
Results for pick and put. We analyse the performance of the P&R model separated by the action (i.e. pick up or put down) in Fig. 4. P&R performs well on both, but pick-up motions are relatively more challenging than put-down actions, especially for priming ability on MoGaze.
@@
Qualitative Results. We demonstrate qualitative examples of our P&R generated motions in Fig. 5. Generated P&R motions appear natural. Starting with an initial pose and velocity, our generated motion first primes the target object (see arrow) and then reaches to it with one of the hands. Evidently, using the goal location matches better the ground-truth. However, using the object location condition solely successfully synthesises reach but positions the body in a different location at the goal. We demonstrate this ability on various target locations including challenging ones, where the location can be located low or high.
@@
Table 3: Impact of condition: We show how each of our modified condition impact P&R model’s performance (last row).
@@
40.14 30.62 45.50 0.38 18.45 10.94 73.21 0.49
38.71 83.44 27.30 0.37 41.07 87.05 28.57 0.46
42.84 88.15 22.41 0.34 36.83 87.57 24.33 0.35
56.76 86.31 16.90 0.31 38.39 94.57 23.59 0.38
56.23 87.66 17.57 0.24 41.00 91.52 23.14 0.33
59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
Table 4: Impact of pre-training. To validate the effectiveness of pre-training on Nymeria, we show the P&R model’s performance trained from scratch or pre-trained on HumanML3D.
@@
No pre-train 49.76 83.86 22.18 0.30 37.20 87.50 24.11 0.33
HumanML3D 56.23 87.56 18.96 0.28 38.39 93.45 24.85 0.34
Nymeria (Ours) 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
@@ 5.5 Ablation and Analysis
We ablate the proposed P&R model our two largest datasets: HD-EPIC and MoGaze. These cover both estimated motion (using EgoAllo for HD-EPIC) and MoCap data (in MoGaze), to ensure our ablation covers both cases.
Condition Ablation. As explained in Sec. 4.2, the proposed P&R method uses text, initial state (initial pose and velocity), and target location as conditions to generate motion. We ablate the impact of each of these conditions in Tab. 3. Using the object location condition gives a significant boost in all metrics, with maximum gains of $82.6\%$ for reach success and $50.3\%$ in location error on MoGaze. This showcases the difficulty of the task, and that it is not trivial to synthesise P&R motions without knowledge of the target. Using the initial state of the body as a condition helps to improve the priming ability of the generated motion, leading to a gain of $20.3\%$ in prime success on HD-EPIC. We find that having both initial pose and initial velocity as initial state conditions is important, especially for the prime success, with drops of at least $2.3\%$ when either is removed on HD-EPIC. Finally, the ablations show that having action knowledge via text (i.e., drop or pick) also improves P&R motion generation on most metrics.
Impact of pre-training. We find that the Nymeria pre-trained model gives a better initialisation (Tab. 4). This is probably because Nymeria has fine-grained motion for everyday activities; therefore, it infuses the knowledge that humans naturally prime and reach objects for everyday actions, which serves better for the downstream task.
@@ 6 Conclusion and Future Work
Humans naturally spot or prime an object before reaching it. Previous motion synthesis benchmarks or methods have failed to explore the role of priming for object reaching. To that end, we curate Prime and Reach (P&R) sequences from five datasets using gaze information and object locations. We propose a P&R motion diffusion model that generates full-body motion using goal pose or target location as a condition, along with initial state and text conditioning. We show that the P&R model can generate better prime and reach motion than previous baselines.
Limitation The current method, in line with similar methods [diomataris2024wandr], does not model hand pose (only body up to wrist). Generating hand motion is an interesting future direction due to its relevance to grasping objects upon reach.
@@
At the University of Bristol, we acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR), funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023]. In particular, we acknowledge the usage of GPU Node hours granted as part of the Sovereign AI Unit call project “Gen Model in Ego‑sensed World” (Aug-Nov 2025) as well as the usage of GPU Node hours granted by AIRR Early Access Project ANON-BYYG-VXU6-M (March-May 2025).
@@
Supplementary Material
Appendix A Qualitative Videos
We include the qualitative video on our website https://masashi-hatano.github.io/prime-and-reach/ showcasing predicted motion sequences from our P&R model over different datasets. For each of the sequences, we provide the goal location in green sphere, our goal-pose conditioned prediction in yellow, and the goal-location conditioned synthesis in brown.
Appendix B Further Details on P&R sequence curation
B.1 Slab Test Method for Priming
The Slab Test Method treats the box as the overlapping volume of three infinite slabs (one for each axis), each bounded by a pair of parallel planes. A visualisation of the intersection checks is shown in Figure S1. The algorithm calculates two key parametric distances along the gaze ray. The first, $t_{\text{near}}$, represents the distance to the last slab plane that the ray enters. It is the furthest entry point, marking the moment the ray is inside all three slabs and thus inside the box. The second, $t_{\text{far}}$, is the distance to the first slab plane that the ray exits. It is the nearest exit point, marking the moment the ray leaves the box volume. A valid intersection occurs if the ray enters the box before it exits, as defined by the condition in Equation 5,
@@ B.1 Slab Test Method for Priming
Figure S1: Visualisation of the Slab Test Method [Majercik2018Voxel] for registering primed object interactions. (Left) Primed by Valid Intersection. (Middle) Primed by Near Miss. (Right) No Priming.
To account for near misses where gaze is directed towards an object but does not intersect its bounding box, we employ a proximity check. First, for a given gaze ray originating at $\mathbf{o}_{\text{cam}}$ with direction $\hat{\mathbf{d}}_{\text{gaze}}$, we find the point on the ray, $\mathbf{p}_{\text{closest}}$, that has the minimum distance to the centre of the object’s 3D bounding box, $\mathbf{b}_{\text{centre}}$. This point is found by projecting the vector from the camera to the box centre onto the gaze ray, as shown in Equation 6.
@@
B.2 Estimating Full Body Pose for P&R Sequences
Building upon the priming data collected previously, we require full-body pose sequences of primed object interactions. Our generation pipeline uses EgoAllo [yi2025egoallo], a method that estimates expressive, full-body human motion from egocentric video and SLAM-based camera poses. The model first converts head pose trajectories into a spatially and temporally invariant representation that encodes relative motion with respect to the ground plane. This representation is used to condition a diffusion-based prior that samples local SMPL-H [MANO:SIGGRAPHASIA:2017] parameters: pose, representing per-joint rotations over time for the full body including hands; shape, encoding time-invariant body proportions such as height and limb length; and contact predictions, indicating per-joint contact with the environment to improve realism. The model is trained on human motion sequences from AMASS [mahmood2019amass], augmented with synthetic egocentric head pose trajectories.
For each interaction, we provide the model with a sequence of video frames and their corresponding camera poses to generate an initial sequence of full-body motions. To enhance the fidelity of hand-object interactions, we calculate the 3D wrist and palm poses from Aria MPS models and provide these to the EgoAllo model to align the generated hands with the wrist and hand locations. This alignment step is crucial; we found that without it, the hands in the generated sequence often remain static and unrealistic. Incorporating these poses yields a more accurate representation of hand orientation in our final motion sequences.
@@ B.2 Estimating Full Body Pose for P&R Sequences
B.3 More Statistics of P&R sequences
We show more detailed statistics on each of our curated datasets in Fig. S2. Concretely, the histograms of body movement, hand movement, and prime gap are shown. Body and hand movement measure the maximum displacement of the body or hands within a P&R motion sequence. Prime gap is the duration between the prime time $t_{p}$ and the pick/put event time $t_{e}$.
Figure S2: Histograms of body movement, hand movement, and prime gap in curated P&R sequences.
B.4 Train-Test Splits
For each dataset, we split the source videos into 70% train -30 % test sets. The P&R sequences curated from these videos were automatically distributed to the corresponding subset. HD-EPIC [perrett2025HD-EPIC] has 156 long videos. We selected 70% (109 videos) for training and the remaining for testing. The curated sequences from the 109 videos were used as train P&R sequences. We perform a similar procedure for MoGaze [kratzer2020mogaze], HOT3D [banerjee2025hot3d], and ADT [pan2023aria]. Zheng et al. [zheng2022gimo] proposed a train-test split for GIMO sequences. We use the same split for our curated P&R sequences. Exact train/test split sizes are given in Tab. S1.

### v2

@@ 5.3 Baselines
To assess the ability of generated motion to replicate humans’ P&R behaviour, we benchmark six methods on our curated datasets: one naive baseline and five previous works (one from the text-to-motion generation and four from location-conditioned human motion generation):
  • Static is a naive baseline that uses the average full-body pose of training data and keeps it static. It showcases the difficulty of the dataset.
  • MDM [84]. We evaluate the checkpoint trained on HumanML3D [24] vs our pre-training on Nymeria [57]. We also fine-tune this model, pre-trained on Nymeria, on our training split using only text conditioning.
  • GMD [39], a guided motion diffusion trained on HumanML3D, is a two-stage motion generation method. The first stage generates a root trajectory that guides full-body motion generation in the second stage.
  • DNO [38] is a framework that treats the pre-trained text-to-motion diffusion model as motion priors and optimises the starting latent noise through backpropagation of task-specific gradients without training for each new task.
  • DartControl [101] autoregressively predicts motion primitives from text and past motion. We adapt its original in-betweening scheme for goal pose and location conditioning, optimising latent noise via DNO for 100 steps.
  • WANDR [15] is an autoregressive c-VAE trained on AMASS and CIRCLE for frame-by-frame motion generation. It utilizes intention features to encode goal location and remaining time.
@@ 5.4 Results
In Tab.˜2, we report results of one model trained on all training sequences from different source datasets. We compare our proposed P&R diffusion model against baselines, reporting results separately on each test set. In Supplementary, we report analogous results where we train on each dataset independently.
@@
The naive static baseline performs poorly on all the metrics for all datasets. Its poor prime success (no priming) and reach success ($12\%$ on average) highlight the difficulty of the task on our curated datasets.
@@
Text-conditioned baselines rely solely on the knowledge of the pick/put action to generate motion and have no information about the target location to prime or reach. We report the performance of three variants of MDM [84]. Text-conditioned baselines perform poorly on all the metrics as they lack sufficient guidance for goal location. Even a fine-tuned MDM model on the P&R sequences has poor performance on prime and reach success on most of the datasets, showing that text is not a sufficient condition for prime and reach.
Conditioning with Goal Pose. We compare our method with three recent controllable motion generation approaches: GMD [39], DNO [38], and DartControl [101], each evaluated both in a zero-shot setting and after fine-tuning on our P&R sequences. In contrast, our goal-pose conditioned P&R model consistently achieves the highest prime and reach success across all five datasets, with gains of up to $18.2\%$ in prime success on GIMO and $25.3\%$ in reach success on MoGaze over the best goal-pose baseline.
@@
Figure 5: Qualitative results on 3 datasets: Ground truth sequence in light green, goal-pose conditioned prediction in translucent yellow, and target location conditioned generation in brown. We show the pose at the initial, prime, and reach timesteps. Prime direction for both ground truth and predictions is shown using arrows, and target object location is shown in sphere.
Conditioning with Object Location. For object location conditioning, we compare with WANDR [15], DNO [38], and DartControl [101], where each model was optimised or trained so that the right wrist of the final frame reaches the target location. Similar to goal pose conditioned baselines, we evaluate these methods in a zero-shot setting and after fine-tuning on our dataset. By construction, latent-noise optimisation methods, used in DNO, DartControl, and P&R  achieve near-perfect reach success across most datasets, since the optimisation objective focuses on the final wrist-target distance. In most cases, P&R outperforms zero-shot and fine-tuned baselines on all metrics. Without guidance on the goal pose, the location error increases for all models. P&R achieves the lowest location error in 4 out of the 5 datasets (2nd best on HOT3D). P&R achieves the best prime success on all datasets.
Impact of fine-tuning on P&R sequences. For both goal-pose and object-location conditioned generation, all baselines generally benefit from fine-tuning on our P&R sequences, resulting in higher prime and reach success on most datasets. In particular, GMD achieves average improvements of $10.2\%$ and $8.1\%$ in prime success and reach success, respectively. This highlights the usefulness of our curated P&R sequences for natural human motion generation.
Why P&R outperforms optimisation-based baselines. Unlike prior methods that mainly use the goal pose or object location as an optimisation target on the final frame, P&R is trained to generate full trajectories conditioned on both the initial state and the goal, encouraging it to learn natural priming strategies (i.e., how to move head, torso, and body to first attend to the future interaction location) rather than merely steering the endpoint. As a result, P&R achieves substantially higher prime success while maintaining strong reach performance.
@@
Results for pick and put. We analyse the performance of the P&R model separated by the action (i.e., pick up or put down) in Fig.˜4. Overall, pick-up motions are relatively more challenging than put-down actions, especially for priming ability on MoGaze when conditioned on the goal pose.
@@
Qualitative Results. We demonstrate qualitative examples of our P&R generated motions in Fig.˜5. Generated P&R motions appear natural. Starting with an initial pose and velocity, our generated motion first primes the target object (see arrow) and then reaches it with one of the hands. Evidently, using the goal location matches better the ground-truth. However, using the object location condition solely successfully synthesises reach but positions the body in a different location at the goal. We demonstrate this ability on various target locations, including challenging ones, where the location can be located low or high.
@@
Table 3: Impact of condition: We show how each of our modified conditions and optimisation impacts P&R model’s performance (last row).
@@
28.50 30.62 65.50 0.45 13.75 10.94 73.21 0.79
39.61 79.64 31.30 0.38 50.67 81.15 49.57 0.69
42.25 81.15 28.41 0.40 52.80 87.57 42.67 0.71
46.90 84.31 29.00 0.38 56.72 91.57 41.59 0.65
48.78 88.46 27.55 0.30 57.34 92.52 37.79 0.60
49.60 88.68 27.30 0.30 59.67 92.64 37.70 0.59
51.00 100.00 25.26 0.27 62.37 100.00 36.43 0.56
Table 4: Impact of pre-training. To validate our pre-training on Nymeria, we show the P&R model’s performance without pre-training and pre-trained on HumanML3D.
@@
No pre-train 41.20 98.86 42.18 0.36 49.50 99.50 48.11 0.68
HumanML3D [24] 49.90 100.00 32.76 0.30 57.45 100.00 40.85 0.59
Nymeria (Ours) 51.00 100.00 25.26 0.27 62.37 100.00 36.43 0.56
@@ 5.5 Ablation and Analysis
We ablate the proposed P&R model on our two largest datasets: HD-EPIC and MoGaze. These cover both estimated motion (using EgoAllo for HD-EPIC) and MoCap data (in MoGaze).
Condition Ablation. As explained in Sec.˜4.1, the proposed P&R method uses text, initial state (pose and velocity), and target location as conditions. We ablate the impact of each of these conditions in Tab.˜3. Using the object location condition gives a significant boost in all metrics, with maximum gains of $81.7\%$ for reach success and $35.5\%$ in location error on MoGaze. This showcases the difficulty of the task, and that it is not plausible to synthesise P&R motions without knowledge of the target. Using the initial state of the body as a condition helps to improve the priming ability of the generated motion, leading to a gain of $10.0\%$ in prime success on HD-EPIC. We find that having both initial pose and initial velocity as initial state conditions is important, especially for the prime success, with drops of at least $2.7\%$ when either is removed on HD-EPIC. Finally, the ablations show that having action knowledge via text (i.e., drop or pick) also improves P&R motion generation on most metrics.
Impact of latent noise optimisation $\mathcal{L}_{\text{opt}}$. As noted in prior work, adopting this optimisation notably improves performance across metrics (except for prime success in MoGaze, which drops marginally).
Impact of pre-training. We find that the Nymeria pre-trained model gives a better initialisation (Tab.˜4) thanks to its diverse, natural human interactions with objects, including intentional interaction motions.
@@ 6 Conclusion and Future Work
Humans naturally spot or prime an object before reaching it. Previous motion synthesis benchmarks or methods have failed to explore the role of priming for object reaching. To that end, we curate Prime and Reach (P&R) sequences from five datasets using gaze information and object locations. We propose a P&R motion diffusion model that generates full-body motion using goal pose or target location as a condition, along with initial state and text conditioning. Results demonstrate improved generation compared to prior baselines.
Limitation Similar to other works [15, 101, 38], we do not model hand pose (only body up to wrist). Generating hand motion is an interesting future direction due to its relevance to grasping objects upon reach.
@@
At the University of Bristol, we acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR), funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023]. In particular, we acknowledge the usage of GPU Node hours granted as part of the AIRR Gateway project “HOI Foundational Model from Egocentric Data” (Dec 2025 - Mar 2026), Sovereign AI Unit call project “Gen Model in Ego‑sensed World” (Aug-Nov 2025) as well as the usage of GPU Node hours granted by AIRR Early Access Project ANON-BYYG-VXU6-M (March-May 2025).
Appendix 0.A Qualitative Videos
We include the qualitative videos on our website https://masashi-hatano.github.io/prime-and-reach/ showcasing predicted motion sequences from our P&R model over different datasets. For each of the sequences, we provide the goal location in green sphere, our goal-pose conditioned prediction in yellow, the goal-location conditioned synthesis in brown and corresponding ground truth motion in green.
Appendix 0.B Further Details on P&R sequence curation
0.B.1 Slab Test Method for Priming
The Slab Test Method treats the box as the overlapping volume of three infinite slabs (one for each axis), each bounded by a pair of parallel planes. A visualisation of the intersection checks is shown in Fig.˜S1. The algorithm calculates two key parametric distances along the gaze ray. The first, $t_{\text{near}}$, represents the distance to the last slab plane that the ray enters. It is the furthest entry point, marking the moment the ray is inside all three slabs and thus inside the box. The second, $t_{\text{far}}$, is the distance to the first slab plane that the ray exits. It is the nearest exit point, marking the moment the ray leaves the box volume. A valid intersection occurs if the ray enters the box before it exits, as defined by the condition in Equation 4,
@@ 0.B.1 Slab Test Method for Priming
Figure S1: Visualisation of the Slab Test Method [59] for registering primed object interactions. (Left) Primed by Valid Intersection. (Middle) Primed by Near Miss. (Right) No Priming.
[…]
@@
0.B.2 Estimating Full Body Pose for P&R Sequences
Building upon the priming data collected previously, we require full-body pose sequences of primed object interactions. Our generation pipeline uses EgoAllo [94], a method that estimates expressive, full-body human motion from egocentric video and SLAM-based camera poses. The model first converts head pose trajectories into a spatially and temporally invariant representation that encodes relative motion with respect to the ground plane. This representation is used to condition a diffusion-based prior that samples local SMPL-H [73] parameters: pose, representing per-joint rotations over time for the full body including hands; shape, encoding time-invariant body proportions such as height and limb length; and contact predictions, indicating per-joint contact with the environment to improve realism. The model is trained on human motion sequences from AMASS [58], augmented with synthetic egocentric head pose trajectories.
For each interaction, we provide the model with a sequence of video frames and their corresponding camera poses to generate an initial sequence of full-body motions. To enhance the fidelity of hand-object interactions, we estimate the 3D wrist and palm poses from Aria MPS models and provide these to the EgoAllo model to align the generated hands with the wrist and hand locations. We found that without this alignment step, the fidelity of the hands in the generated sequence is often diminished. Incorporating these poses yields a more accurate representation of hand positions and their orientations in our final motion sequences.
@@ 0.B.2 Estimating Full Body Pose for P&R Sequences
0.B.3 Comparing EgoAllo to Mocap.
To verify the suitability of EgoAllo outputs as an approximation to the standard 3D body pose annotations, typically acquired using Mocap, we compare the results of P&R when using EgoAllo in place of the Mocap data on the ADT [62] dataset. We chose this dataset as it uses the Aria glasses, making it suitable for EgoAllo body pose estimations. One thing to note is that hand (palm and wrist) tracking data is not available from the Aria glasses in this dataset, which are known to improve the EgoAllo estimations.
The results are shown in Tab.˜S1. We provide results when training using Mocap and EgoAllo as well as when evaluating on the test set poses from Mocap (left) and EgoAllo (right). Overall, P&R trained on EgoAllo body pose estimates performs slightly worse than the corresponding ground truth. When conditioning with object location, EgoAllo performs similarly to Mocap (e.g. 97.64% vs. 100% Reach Success). While there is a performance drop in Reach Success under the goal pose condition when evaluating on the Mocap test set (51.76% vs. 70.58%), the broader motion-based metrics (e.g. Goal MPJPE, MPJPE, Loc Err and Foot Skating) still remain comparable. For example, MPJPE drops by <0.1m<0.1m in every case when comparing models trained on EgoAllo to those trained on Mocap. This indicates that the overall motion quality is acceptable.
Table S1: Comparison of EgoAllo outputs and ground truth. We compare our P&R method on ADT [62] when trained and evaluated on both ground-truth Mocap data and EgoAllo.
ADT (Mocap) ADT (EgoAllo)
Condition
Training
Data
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
+ Initial State & Goal Pose EgoAllo 29.41 51.76 0.17 5.88 0.38 0.17 28.05 81.17 0.18 4.87 0.42 0.18
Mocap 37.64 70.58 0.08 1.17 0.33 0.11 31.70 92.68 0.13 1.22 0.38 0.21
+ Initial State & Object Loc. EgoAllo 45.88 97.64 0.55 52.94 0.54 0.18 36.58 96.34 0.53 45.12 0.54 0.18
Mocap 43.53 100.00 0.55 52.54 0.48 0.11 39.02 100.00 0.54 45.12 0.51 0.20
0.B.4 More Statistics of P&R sequences
We show more detailed statistics on each of our curated datasets in Fig.˜S2. Concretely, the histograms of body movement, hand movement, and prime gap are shown. Body and hand movement measure the maximum displacement of the body or hands within a P&R motion sequence. Prime gap is the duration between the prime time $t_{p}$ and the pick/put event time $t_{e}$.
Figure S2: Statistics of the curated P&R sequences. In a, b, and c, we highlight the high diversity of our goal locations by plotting the histogram of goal distance, elevation, and azimuthal location relative to the start location. In d, we plot the histogram of body movement in our sequences. In e, we plot how the gaze shifts from the goal during the prime and reach motion, where the shift is minimum when the gaze is on the object. Finally in f, we plot the histogram for time interval between priming and reaching.
0.B.5 Train-Test Splits
For each dataset, we split the source videos into 70% train -30 % test sets. The P&R sequences curated from these videos were automatically distributed to the corresponding subset. HD-EPIC [67] has 156 long videos. We selected 70% (109 videos) for training and the remaining for testing. The curated sequences from the 109 videos were used as train P&R sequences. We perform a similar procedure for MoGaze [42], HOT3D [6], and ADT [62]. Zheng et al. [102] proposed a train-test split for GIMO sequences. We use the same split for our curated P&R sequences. Exact train/test split sizes are given in Tab.˜S2.